OpenStudy (anonymous):

Two cars travel between towns which are 50 miles apart. Both cars start from the same town at the same time. The first car travels at 50 mph while the second travels at 30 mph. If the first car stops in the second town for 5 minutes and then returns (still at 50 mph), determine where the two cars will meet.

OpenStudy (phi):

The first car takes 1 hour to arrive at the second town. It waits an additional 5/60 = 1/12 of an hour. total time 13/12 of an hour the second car has traveled r* t= 30 * 13/12 = 32.5 miles. It is 50-32.5 = 17.5 miles from the second city

OpenStudy (phi):

so we now have this problem